# Adaptive Kernel Selection for Stein Variational Gradient Descent

### v1

@@ 1 Introduction
The performance of SVGD depends critically on the choice of the kernel function, as it determines the interaction between particles and the overall convergence of the method (see also Figures 6 and 1). Convergence results typically refer to mean-field convergence with respect to KSD, whose relation to weak convergence depends on the selected kernel [10]. The commonly used median heuristic [11] provides a simple implementation but lacks theoretical justification and is known to degrade in performance as the dimensionality of the problem increases [2, 34, 32]. Recent work has developed tools to mitigate performance degradation in high-dimensional settings [6, 9, 18].
@@
    Theoretical Analysis. We provide theoretical motivation by analyzing our algorithm in the discrete-time mean-field setting, extending existing convergence results for SVGD with fixed kernels. Specifically, we show that the supremum of the KSD over a parameterized kernel class converges to zero as the particle distribution approaches the target.
@@ 2 Mathematical background
We begin by considering a symmetric positive definite kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ and its associated RKHS $\mathcal{H}_{0}$. We define $\mathcal{H}$ as the $d$-fold Cartesian product of $\mathcal{H}_{0}$ equipped with the inner product $\langle f,g\rangle_{\mathcal{H}}=\sum_{i=1}^{d}\langle f_{i},g_{i}\rangle_{\mathcal{H}_{0}}$ and the canonical feature map $\Phi_{k}(x)=k(\cdot,x)\in\mathcal{H}_{0}$. Moreover, we denote by $x\cdot y$ the Euclidean inner product and $\nabla\cdot$ the divergence operator. The space of probability measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d})$ is denoted by $\mathcal{P}(\mathbb{R}^{d})$ and $\mathcal{P}_{p}(\mathbb{R}^{d})$ denotes the subspace of measures with finite $p$-th moment. For $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, we define the Wasserstein $p$-distance
$\mathcal{W}_{p}(\mu,\nu)=\inf_{\gamma\in\Gamma(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\lVert x-y\rVert_{2}^{p}\mathrm{d}\gamma(x,y),$
@@
where $B(\mathcal{H})$ denotes the unit ball in $\mathcal{H}$. This optimization problem is solved by $f^{\ast}=\frac{\psi}{\lVert\psi\rVert_{\mathcal{H}}}$ with $\psi=\int_{\mathbb{R}^{d}}\mathcal{A}_{\pi}^{k}\mathrm{d}\mu$, where $\mathcal{A}_{\pi}^{k}(x)=\nabla\log\pi(x),\Phi_{k}(x)+\nabla\Phi_{k}(x)\in\mathcal{H}$, and $\Phi_{k}$ is the feature map associated with the kernel $k$. As a result, the supremum evaluates to the RKHS norm of $\psi$, giving the equivalent characterization $\mathrm{KSD_{#3}}(\mu|\pi)=\lVert\psi\rVert_{\mathcal{H}}$.
@@
Given a target distribution $\pi$ and reference distribution $\mu_{0}$, SVGD transforms $\mu_{0}$ into an approximation of $\pi$ by choosing $\mu_{n+1}\coloneqq T_{\sharp}\mu_{n}$, where $T_{\sharp}\cdot$ is the push-forward operator for the map $T(x)=x+\gamma\psi^{\mu_{n}}(x)$, with the vector field $\psi^{\mu}=\int_{\mathbb{R}^{d}}\mathcal{A}_{\pi}^{k}\mathrm{d}\mu$ for $\mu\in\mathcal{P}(\mathbb{R}^{d})$ being the direction of steepest descent. This is motivated by the fact that the solution of Equation 1 implies that $\frac{\psi^{\mu_{n}}}{\lVert\psi^{\mu_{n}}\rVert_{\mathcal{H}}}$ is the minimizer of $\left.\frac{\mathrm{d}}{\mathrm{d}\gamma}\mathrm{KL}(T_{\sharp}\mu_{n}\,\|\,\pi)\right|_{\gamma=0}$ in the unit ball of $\mathcal{H}$ [[, cf.]Theorem 3.1]SVGD, where $\mathrm{KL}(\cdot\,\|\,\cdot)$ denotes the KL-divergence. In particular, for this choice we have
$\frac{\mathrm{d}}{\mathrm{d}\gamma}\mathrm{KL}\left(\big(\operatorname{Id}+\gamma\psi^{\mu_{n}}\big)_{\sharp}\mu_{n}\,\middle\|\,\pi\right)\Big|_{\gamma=0}=-\mathrm{KSD_{#3}}(\mu_{n}|\pi).$ (2)
@@ 3 Adaptive kernel selection for SVGD
We implement this idea by adjusting the kernel parameter using one or possibly more gradient ascent steps with a step size $s>0$ for KSD before executing the particle update and also allow the possibility of not updating the parameter at every step of SVGD to decrease runtime (see Algorithm 1). It is worth noting that the base SVGD step for transporting the particles can be replaced by more advanced variants (e.g., adaptive step-size schedules, line searches, momentum, or second-order/preconditioned updates) without modifying our kernel-selection mechanism. Similarly, we may replace the gradient ascent on the kernel parameter with alternative optimization schemes.
Our implementation of the kernel update is based on the formula [5, 16]
@@ 4.1 Convergence analysis of SVGD under Talagrand’s inequality
$\mathcal{F}(\tilde{\mu})\leq\mathcal{F}(\mu)+\gamma\langle\psi_{\theta}^{\mu},g\rangle_{\mathcal{H}_{\theta}}+\frac{\gamma^{2}K}{2}\|g\|_{\mathcal{H}^{\theta}}$ (6)
with $K=(\alpha^{2}+L)B$. For the iterates (3) with a fixed kernel parameter $\theta\in\Theta$ one can then derive the descent condition [28, Theorem 3.2],
@@
The key aspect to verify (7) is the verification of $\gamma\|\psi_{\theta}^{\mu_{n}}\|_{\mathcal{H}^{\theta}}\leq\frac{\alpha-1}{\alpha B}$ using Talagrand’s inequality, Assumption 2, which allows to apply (6) for $g=\psi_{\theta}^{\mu_{n}}$. Condition (7) can then be used to argue that
@@
$k_{\theta}(x,y)=\sum_{n=1}^{N}\theta_{i}k_{i}(x,y)\quad\text{with}\ \sum_{n=1}^{N}\theta_{i}=1\,.$
For the general setting, the kernel parameter update $\theta_{n}$ in (9) needs to be approximated. Note that in the general setup we assume that $\operatorname*{argmax}_{\theta\in\Theta}\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi))\neq\emptyset$ for all $n\in\mathbb{N}$.
@@
Under the same assumptions as in Lemma 1 it holds that
$\lim_{n\to\infty}\max_{\theta\in\Theta}\ \mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)=0\,.$
@@
We assume that there exists a sequence $(\varepsilon_{n})_{n\in\mathbb{N}}$ such that $\sum_{n=0}^{\infty}\varepsilon_{n}<\infty$ and
$\max_{\theta\in\Theta}{\mathrm{KSD}}_{\theta}^{2}(\mu_{n}\mid\pi)-{\mathrm{KSD}}_{\theta_{n+1}}^{2}(\mu_{n}\mid\pi)\leq\varepsilon_{n}$
for all $n\in\mathbb{N}$.
@@
it holds that
$\lim_{n\to\infty}\max_{\theta\in\Theta}\mathrm{KSD_{#3}}(\mu_{n}|\pi)=0\,$,
where $(\mu_{n})_{n\in\mathbb{N}}$ is generated by (10).
@@
$=\mathcal{F}(\mu_{n})-c_{\gamma}\max_{\theta\in\Theta}\,\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)+c_{\gamma}\big(\max_{\theta\in\Theta}\,\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)-\,\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)\big)$
$\leq\mathcal{F}(\mu_{n})-c_{\gamma}\max_{\theta\in\Theta}\,\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)+c_{\gamma}\varepsilon_{n}\,.$
Iterating this inequality over $n$ yields
$c_{\gamma}\sum_{n=0}^{\infty}\max_{\theta\in\Theta}\,\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)\leq\mathcal{F}(\mu_{0})+c_{\gamma}\sum_{n=0}^{\infty}\varepsilon_{n}<\infty\,.$
@@ 5.1 Kernel parameterization
where $\lVert\cdot\rVert_{p}$ denotes the $p$-norm on $\mathbb{R}^{d}$ [[, e.g.]]SVGD,DBLP:conf/iclr/BaEGSSWZ22,Duncan2019OnTG. We will focus on selection strategies for the parameter $h$, which is known as the kernel bandwidth. The commonly used heuristic sets $h=\frac{\text{med}^{p}}{\log(M-1)}$, where med denotes the current median distance between the particles. This choice is motivated by the goal of balancing the two terms in the SVGD update (3[17].
@@ 5.3 Linear inverse problem based on ODE
The following example is adapted from [33, Example 2.1]. We consider the one-dimensional differential equation
@@
For the Bayesian formulation of the inverse problem, we consider a Gaussian process (GP) prior for $u(\cdot)$ given by the truncated Karhunen-Loève (KL) expansion
@@ A.2 Scaling dimension: multivariate normal distribution
Lastly, we focus on the marginal particle distributions for $M=200$ and $d=8$. To ease the visualization, we normalized them by scaling each component of the particles with the inverse of the corresponding marginal standard deviation of the target distribution (i.e. we multiplied the $i$-th component with $i$). This turns each marginal distribution of $\pi_{d}$ into a standard normal distribution. Figures 8 and 9 show the histograms of these normalized marginal particles distributions for $d=8$ generated using Med-SVGD and Ad-SVGD, respectively. A standard Gaussian density is shown in each plot for comparison. Again, we observe that Ad-SVGD is able to capture all marginal distributions well, while Med-SVGD underestimates the uncertainty of the target distribution. These observations are also visible in Figures 10 and 11, where the quantiles of the normalized marginal particle distributions are compared against the target quantiles (i.e. against a standard normal distribution).

### v2

@@ 1 Introduction
The performance of SVGD depends critically on the choice of the kernel function, as it determines the interaction between particles and the overall convergence of the method (see also Figures 9 and 1). Convergence results typically refer to mean-field convergence with respect to KSD, whose relation to weak convergence depends on the selected kernel [11]. The commonly used median heuristic [12] provides a simple implementation but lacks theoretical justification and is known to degrade in performance as the dimensionality of the problem increases [2, 40, 36]. Recent work has developed tools to mitigate performance degradation in high-dimensional settings [7, 37, 10, 21]. The algorithm has also been modified through the use of neural networks to learn the update direction [17, 39].
@@
In this work, we address the fundamental research question of whether SVGD, without any architectural or algorithmic enhancements, can achieve competitive performance when equipped with a principled and adaptively chosen kernel. By isolating kernel selection from other modifications found in SVGD variants, we aim to rigorously understand the extent to which kernel choice alone governs the effectiveness of SVGD and how far adaptive kernel learning can push the capabilities of the original method.
@@
    Theoretical Analysis. We provide theoretical motivation by analyzing our algorithm in the discrete-time mean-field setting, extending existing convergence results for SVGD with fixed kernels. Specifically, we show that the supremum of the KSD over a parameterized kernel class converges to zero as the particle distribution approaches the target. Assuming a Stein logarithmic Sobolev inequality we further derive iteration complexity in the mean-field limit.
@@ 2 Mathematical background
We begin by considering a symmetric positive definite kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ and its associated RKHS $\mathcal{H}_{0}$. We define $\mathcal{H}$ as the $d$-fold Cartesian product of $\mathcal{H}_{0}$ equipped with the inner product $\langle f,g\rangle_{\mathcal{H}}=\sum_{i=1}^{d}\langle f_{i},g_{i}\rangle_{\mathcal{H}_{0}}$ and the canonical feature map $\Phi_{k}(x)=k(\cdot,x)\in\mathcal{H}_{0}$. Moreover, we denote by $x\cdot y$ the Euclidean inner product and $\nabla\cdot$ the divergence operator. The space of probability measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ is denoted by $\mathcal{P}(\mathbb{R}^{d})$ and $\mathcal{P}_{p}(\mathbb{R}^{d})$ denotes the subspace of measures with finite $p$-th moment. For $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, we define the Wasserstein $p$-distance
$\mathcal{W}_{p}(\mu,\nu)=\inf_{\gamma\in\Gamma(\mu,\nu)}\Big(\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\lVert x-y\rVert_{2}^{p}\mathrm{d}\gamma(x,y)\Big)^{1/p},$
@@
where $B(\mathcal{H})$ denotes the unit ball in $\mathcal{H}$. This optimization problem is solved by $f^{\ast}=\frac{\psi}{\lVert\psi\rVert_{\mathcal{H}}}$ with $\psi=\int_{\mathbb{R}^{d}}\mathcal{A}_{\pi}^{k}\mathrm{d}\mu$, where $\mathcal{A}_{\pi}^{k}(x)=\nabla\log\pi(x)\,\Phi_{k}(x)+\nabla\Phi_{k}(x)\in\mathcal{H}$, and $\Phi_{k}$ is the feature map associated with the kernel $k$. As a result, the supremum evaluates to the RKHS norm of $\psi$, giving the equivalent characterization $\mathrm{KSD_{#3}}(\mu|\pi)=\lVert\psi\rVert_{\mathcal{H}}$.
@@
Given a target distribution $\pi$ and reference distribution $\mu_{0}$, SVGD transforms $\mu_{0}$ into an approximation of $\pi$ by choosing $\mu_{n+1}\coloneqq T_{\sharp}\mu_{n}$, where $T_{\sharp}\cdot$ is the push-forward operator for the map $T(x)=x+\gamma\psi^{\mu_{n}}(x)$, with the vector field $\psi^{\mu}=\int_{\mathbb{R}^{d}}\mathcal{A}_{\pi}^{k}\mathrm{d}\mu$ for $\mu\in\mathcal{P}(\mathbb{R}^{d})$ being the direction of steepest descent. This is motivated by the fact that the solution of Equation 1 implies that $\frac{\psi^{\mu_{n}}}{\lVert\psi^{\mu_{n}}\rVert_{\mathcal{H}}}$ is the minimizer of $\left.\frac{\mathrm{d}}{\mathrm{d}\gamma}\mathrm{KL}(T_{\sharp}\mu_{n}\,\|\,\pi)\right|_{\gamma=0}$ in the unit ball of $\mathcal{H}$ [18, cf.], where $\mathrm{KL}(\cdot\,\|\,\cdot)$ denotes the KL-divergence. In particular, for this choice we have
$\frac{\mathrm{d}}{\mathrm{d}\gamma}\mathrm{KL}\left(\big(\operatorname{Id}+\gamma\psi^{\mu_{n}}\big)_{\sharp}\mu_{n}\,\middle\|\,\pi\right)\Big|_{\gamma=0}=-\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi).$ (2)
@@ 3 Adaptive kernel selection for SVGD
We implement this idea by adjusting the kernel parameter using one or possibly more gradient ascent steps with a step size $s>0$ for KSD before executing the particle update and also allow the possibility of not updating the parameter at every step of SVGD to decrease runtime (see Algorithm 1). To enable this option, we introduce a user-specified decision variable $\mathrm{paramupdate}$. It is worth noting that the base SVGD step for transporting the particles can be replaced by more advanced variants (e.g., adaptive step-size schedules, line searches, momentum, or second-order/preconditioned updates) without modifying our kernel-selection mechanism. Similarly, we may replace the gradient ascent on the kernel parameter with alternative optimization schemes. Our implementation of the kernel update is based on the formula [6, 20]
@@
   Decide $\mathrm{paramupdate}\in\{\mathrm{True},\mathrm{False}\}$;
@@ 4.1 Convergence analysis of SVGD under Talagrand’s inequality
$\mathcal{F}(\tilde{\mu})\leq\mathcal{F}(\mu)+\gamma\langle\psi_{\theta}^{\mu},g\rangle_{\mathcal{H}_{\theta}}+\frac{\gamma^{2}K}{2}\|g\|_{\mathcal{H}_{\theta}}$ (6)
with $K=(\alpha^{2}+L)B$. For the iterates (3) with a fixed kernel parameter $\theta\in\Theta$ one can then derive the descent condition [32, Theorem 3.2],
@@
The key aspect to verify (7) is the verification of $\gamma\|\psi_{\theta}^{\mu_{n}}\|_{\mathcal{H}_{\theta}}\leq\frac{\alpha-1}{\alpha B}$ using Talagrand’s inequality, Assumption 2, which allows to apply (6) for $g=\psi_{\theta}^{\mu_{n}}$. Condition (7) can then be used to argue that
@@
$k_{\theta}(x,y)=\sum_{i=1}^{N}\theta_{i}k_{i}(x,y)\quad\text{with}\ \sum_{i=1}^{N}\theta_{i}=1\,.$
For the general setting, the kernel parameter update $\theta_{n}$ in (9) needs to be approximated. Note that in the general setup we assume that $\operatorname*{argmax}_{\theta\in\Theta}\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)\neq\emptyset$ for all $n\in\mathbb{N}$.
@@
Under the same assumptions as in Lemma 1 it holds that $\lim_{n\to\infty}\max_{\theta\in\Theta}\ \mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)=0\,.$
@@
We assume that there exists a sequence $(\varepsilon_{n})_{n\in\mathbb{N}}$ such that $\lim_{n\to\infty}\varepsilon_{n}=0$ and
$\max_{\theta\in\Theta}{\mathrm{KSD}}_{\theta}^{2}(\mu_{n}\mid\pi)-{\mathrm{KSD}}_{\theta_{n}}^{2}(\mu_{n}\mid\pi)\leq\varepsilon_{n}\qquad\text{for all }n\in\mathbb{N}.$
@@
it holds that $\lim_{n\to\infty}\max_{\theta\in\Theta}\mathrm{KSD_{#3}}(\mu_{n}|\pi)=0\,$, where $(\mu_{n})_{n\in\mathbb{N}}$ is generated by (10).
@@
Since $(\mathcal{F}(\mu_{n}))_{n\in\mathbb{N}}$ is decreasing and bounded below by zero, it converges and, in particular, it is a Cauchy sequence. Hence, we have
$c_{\gamma}\,\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)\leq F(\mu_{n})-F(\mu_{n+1})\to 0$
as $n\to\infty$. Under Assumption 4 we obtain
$\max_{\theta\in\Theta}\,\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)\leq\mathrm{KSD_{#3}^{2}}(\mu_{n}|\pi)+\varepsilon_{n}\to 0$
as $n\to\infty$. ∎
Under the so-called Stein logarithmic Sobolev inequality, we can further extend the convergence of Ad-SVGD to convergence in the KL divergence. Under specific scenarios for $\varepsilon_{n}$ we derive iteration complexity bounds for the mean-field limiting system (10). Moreover, this setting allows us to explicitly describe the bias of Ad-SVGD when the KSD maximization is only solved up to a fixed $\bar{\varepsilon}$-accuracy in each iteration. The details can be found in Appendix B. Finally, we emphasize that we require only one kernel in our kernel class $\{k_{\theta}\}$ to satisfy the Stein logarithmic Sobolev inequality which yields a promising flexibility to verify this condition.
@@ 5 Numerical Experiments
In the following section, we evaluate Ad-SVGD on two numerical experiments: a one-dimensional Gaussian mixture model and a linear inverse problem governed by an ordinary differential equation. Additional toy experiments examining the effect of increasing dimensionality are provided in Appendix A. Moreover, in Section 5.4 we demonstrate the advantages of Ad-SVGD over SVGD with the median heuristic on a Bayesian logistic regression model, which was also used in prior SVGD literature.
@@ 5.1 Kernel parameterization
where $\lVert\cdot\rVert_{p}$ denotes the $p$-norm on $\mathbb{R}^{d}$ [18, 2, 8, e.g.]. We will focus on selection strategies for the parameter $h$, which is known as the kernel bandwidth. The commonly used heuristic sets $h=\frac{\text{med}^{p}}{\log(M-1)}$, where med denotes the current median distance between the particles. This choice is motivated by the goal of balancing the two terms in the SVGD update (3[18].
@@
Beyond our current experiments, the flexibility of Ad-SVGD allows for more scalable kernel families in higher-dimensional settings. In particular, spectral or low-rank kernel parametrizations can adapt the dominant directions of variability without incurring the exponential cost of full product kernels, while Matérn or mixed-product constructions provide additional control over smoothness and anisotropy.
@@ 5.3 Linear inverse problem based on ODE
The following example is adapted from [38, Example 2.1]. We consider the one-dimensional differential equation
@@
For the Bayesian formulation of the inverse problem, we consider a Gaussian process (GP) prior for $u$ given by the truncated Karhunen-Loève (KL) expansion
@@
5.4 Bayesian logistic regression
As in [18] and [21], we consider a Bayesian logistic regression (BLR) model applied to the Covertype data set [4]. For the regression weights $w$, we assign a Gaussian prior $w\mid\alpha\sim\mathcal{N}(0,\alpha^{-1})$ with $\alpha\sim\mathrm{Gamma}(1,0.01)$ and we want to infer the posterior of $x=[w,\alpha]$. Following [21], we use the No-U-Turn Sampler (NUTS) introduced by [13] to generate a reference for evaluating the posterior approximation quality of Med-SVGD and Ad-SVGD.
Figure 5: Approximation quality of SVGD particles for Bayesian logistic regression measured by prediction accuracy (left) and $\mathrm{MMD}^{2}$ to reference samples (right)
Again following [21], we test the methods on subsets of the original data set of size 1000 and aggregate our results over 10 random draws. To evaluate the performance of the methods for the classification task, we compute the prediction accuracy over a test set achieved using the particle mean. As shown in Figure 5, Ad-SVGD outperforms Med-SVGD across almost all seeds, achieving a higher test accuracy while also being closer to the MCMC reference.
Figure 6: Bandwidth evolution for BLR example
To measure the accuracy of the posterior approximation, the same figure also shows the squared maximum mean discrepancy ($\mathrm{MMD}^{2}$) [25, Section 3.5] to the MCMC samples. We observe a significant difference of how well the two methods are able to capture the posterior distribution. This is further confirmed by Figure 7, which shows the covariance matrices of the final particles compared to the MCMC reference, averaged over the 10 seeds (this visualization style is taken from [21]). We observe that Med-SVGD severely underestimates the posterior uncertainty, while Ad-SVGD gives a better approximation. Additionally, Figure 6 shows the evolution of the bandwidths for each component. We see that from an uninformed initialization (1 in every component), Ad-SVGD is able to recover suitable bandwidths, which differ from the initialization by several orders of magnitude.
Figure 7: Covariance matrices of final SVGD particle sets compared to MCMC samples (mean over 10 seeds)
@@ A.2 Scaling dimension: multivariate normal distribution
Lastly, we focus on the marginal particle distributions for $M=200$ and $d=8$. To ease the visualization, we normalized them by scaling each component of the particles with the inverse of the corresponding marginal standard deviation of the target distribution (i.e. we multiplied the $i$-th component with $i$). This turns each marginal distribution of $\pi_{d}$ into a standard normal distribution. Figures 11 and 12 show the histograms of these normalized marginal particles distributions for $d=8$ generated using Med-SVGD and Ad-SVGD, respectively. A standard Gaussian density is shown in each plot for comparison. Again, we observe that Ad-SVGD is able to capture all marginal distributions well, while Med-SVGD underestimates the uncertainty of the target distribution. These observations are also visible in Figures 13 and 14, where the quantiles of the normalized marginal particle distributions are compared against the target quantiles (i.e. against a standard normal distribution).